To: J. Albertine, D. Helms, W. Skinner
cc: Division (M7), FEL Coordination Group
From: F. Dylla
Subject: IRFEL Weekly Report, June 16-20, 1997
Date: June 20, 1997
Highlights for the week: we are on track for completing the list
of final activities for the injector test stand that would enable
shut down of the ITS next Monday (June 23). The activity list
included completion of longitudinal emittance and bench length
measurements and an extended run to monitor the dark current in
the gun. On Monday of this week, the 1/4 cryounit was removed
from the ITS and placed in the injector area of the FEL Building.
Removal and transfer of the gun and other ITS hardware slated
for transfer to FEL Building is scheduled to begin on June 23
(see ITS report for details).
This week Jefferson Laboratory hosted a visit from DOE Office
of Energy Research senior management for their review of the laboratory's
institutional plan. The visit included a status report on the
IR Demo project and a tour of the FEL Facility.
Final preparations were made for next week's Laser Processing
Consortium Workshop (June 25-26). Approximately 50 attendees
from outside the laboratory are expected. The focus of the meeting
will be the generation of proposals for initial fit-up and use
of the FEL user labs.
At the request of Navy program office, revised proposals were
forwarded on June 16 for use of the $3.7M of FY97 DOD appropriations
planned for allocation to Jefferson Lab for July 1997-Sept. 1998
Injector Test Stand
Regarding the photocathode gun:
We worked on taking longitudinal measurements at high bunch charges
and followed that activity with a 48-hour run to monitor dark
current versus time and thereby gain some additional information
regarding the field-emission behavior of the gun.
The status of the gun at this writing (0930, 20 Jun 97) is as
follows: We have measurements of bunch length and longitudinal
emittance at bunch charges of 10, 60, and 135 pC. One qualitative
feature observed at high bunch charges is a hollow longitudinal
density profile, one that shows clumping at the head and tail
of the bunch. This observation correlates with simulations (PARMELA),
at least qualitatively, and therefore is attributed to space charge.
However, there remains a question concerning the sensitivity
of the density profile with respect to the solenoid setting.
The drive laser introduced a complication in that it was unstable
during the high-bunch-charge measurements, and this generated
uncertainties in some of the data. For example, at 135 pC, there
was considerable jitter in the energy spread. Moreover, there
was a discrepancy between measurements of quantum efficiency and
output current. This lead to an uncertainty in bunch-charge measurements
at 60 pC; they could in fact have been 40 pC. However, after
revamping the drive laser as described below, and during a subsequent
set of measurements at 10 pC, the drive laser operated very stably.
In view of these circumstances, and to complete a solid set of
experiments, we are embarking on two more 10-hour runs through
the weekend, during which we will repeat and augment the 60 and
135 pC measurements. We made a new cathode yesterday after completing
the field-emission run, and are gearing up to begin the additional
runs today. The plan remains to begin dismantling the gun on
Monday, 23 Jun 97, to prepare for moving it to the FEL Facility,
but this precise date remains tentative pending the success of
this weekend's activities.
The pump volute on the drive laser was replaced last Friday with
one made of PVC. This will eliminate problems with lowered (and
often unstable) power that occurs when using the stock electroless
nickel-coated volutes. The lamps were also cleaned and the deionizer
cartridge replaced. The laser now performs at the level it did
several months ago. At 1700 today (20 Jun 97) it will have operated
this week for 100 hours in support of gun experiments. The power
output was quite stable over this period (+/- 0.1 W at 19.6 W
infrared output). The intensity "ringing" of long macropulses
continues to be investigated. The vendor is no longer convinced
it is due to piezoelectric resonances in the KD*P crystals, but
instead may be in the electronics. We will continue to investigate
this as soon as the laser is not being used for gun experiments.
We are also preparing to move the drive laser as soon as the
utilities are available in the new clean room located in the FEL
Regarding the cryounit:
The cryounit was moved to the FEL Facility this week. Rough alignment
was completed and final alignment started.
Regarding development of 50 kW ceramic warm windows:
We tested a Wesgo ceramic to 50 kW in the resonant ring test assembly
and found very little heating. One of our staff will travel to
SLAC next week to make RF measurements with the goal of optimizing
the window design for low losses. Eight Wesgo ceramics were
sent out to be metallized and are expected back within a week.
Regarding the beam-transport system:
In anticipation of the assembly of the optical chicane dipoles,
Everson Electric completed two trials of the gluing of shim materials
to sample cores. The first, using 30 psi, failed in that the
glue thickness ranged from .0005 inch to .0025 inch, with one
reading at .0045 inch. However, the second trial, at 120 psi,
was successful, with a combined, uniform thickness of two layers
of epoxy at 16 points of less than .0005 inches.
Cores for the reverse-bend dipoles are in final machining, and
the mechanical quality of their cut-to-size mu metal passed qualification.
However, because of the changed dimension of the thickness of
the mu metal and brass associated with the non-interchangeable
parts assembly plan, the before-final-grinding back-leg height
is already at the nominal final dimension. The existing plan
was to grind an additional .010 in. to achieve the final, accurate
dimension which leaves the back legs short by this amount. At
week's end we are working the issue. A simple solution is to
reduce the 180-degree bend's gap height by .010 in. to maintain
the original match between families. Reduction in the size of
the gap affects the chambers and coils. The chamber slot reduction
has been coordinated and cleared with the designer, and the coils
have to have thinner bedding strips for both styles of magnet.
Core parts for the injection/extraction dipoles lack only the
final delivery of mu metal. The manufacturing process has been
qualified, and those parts need only go through their laser cut.
A final detail is being worked out on the position of the bus
leads on the coils before the complete set can be delivered.
Additional parts and modifications to our Magnet Test Stand that
will allow comparative measurement of the dipole magnets are nearing
Sign-off of the Panofsky trim quad drawing set is being delayed
by checking of the assembly tooling. The tooling is being included
with the order because it is necessary for an accurate assembly
by the vendor. Sextupole cores and coils remain on track for
a delivery on 8 Jul 97.
Regarding corrector magnets, the mu-metal phasing dipole is being
drawn up. (Note: We still have to qualify these mu-metal correctors
with regard to cross talk and fringe field.) Three-dimensional
magnet modeling of correctors is continuing, and we seem to be
converging on a design for a constant-perimeter, two-shell air-core-coil
set that will achieve the 1% uniformity specification. This design
could be made by a simple process involving fabrication of a flat
coil followed by bending it around a tube. In parallel, Advanced
Magnet Lab in Florida submitted a preproposal wherein they could
make similar style coils with their unique conductor lay-up process
that would achieve the same quality. However, because of development
and prototyping, they are way out in schedule, and cost is considerably
more than we project for our process.
Regarding vacuum chambers, completion of the injection X-chamber
was delayed as we assessed the impact of the contamination that
our Machine Shop introduced by using cutting oil when machining
the ends. This was contrary to the informal, written procedure
worked up with the Shop to produce the chamber. We are working
on a procedure that will be signed off and which has a higher
probability of being followed. At the same time, inspection of
the weld of the extraction Y-chamber (as well as of all future
chambers) with a boroscope should increase the welder's skill
at producing good welds. The parts for the optical-chicane chambers
are being prepared for weld. We are working to ensure the necessary
manpower is available to complete all of the vacuum-chamber welds
Drawings for the insertable dumps are being signed off, and the
design of the beam scraper is nearing completion.
Regarding cryomodule fabrication:
The status of the cryomodule is as follows: First cryounit --
complete. Second cryounit -- will be moved to cryomodule staging
area next week. Third cryounit -- inserted into the helium vessel;
welding will be completed early next week. Fourth cryounit --
spacer flanges were fabricated and assembled.
Regarding electron-beam instrumentation and controls (I&C):
Six sets of beam instrumentation controls drawings were signed
off. All beam viewer and motor control interface chassis are
The PSS systems are proceeding well, and the ODH system will be
operational by 6/27/97. This is required prior to cooling down
the quarter cryomodule. A review of the MPS is scheduled for
A preliminary design for an analog monitoring system is complete.
A cost comparison is being made between an HP 4:1 mux (24 required
~ $22K) and home built one.
All VME crates are installed, the cryogenic controls are operational.
This IOC is connected to the Central Helium Liquifier directly
by a fiberoptic link. The controls database is being loaded and
instrumentation checkout is proceeding for the July 1 cooldown
of the quarter cryomodule.
Further measurements on the quad girders resulted in a few other
design errors being discovered. The girders must be reworked
before further alignment work can proceed. We are also working
on making up borescoping fixtures to check for potential systematic
errors in the quadrupole alignment which might lead to an offset
of up to 200 microns in the alignment of the quadrupoles and the
This has been a busy week in terms of receipt of optics and components.
We received the sapphire mirrors (coated for 3 micron operation)
from Research Electro-Optics. Their metrology shows the ROC to
be within our spec (+/- 0.2%) for the output couplers, and somewhat
out of spec (+0.4%) for the high reflectors. We will perform
our own checks, and await results from China Lake before determining
if they need to be sent back for refiguring. We also received
a partial order (8 of 18) of the optical transport mirrors from
SPAWAR. We should receive the rest of the order next week. We
received the mirror can stands on Monday, but rejected some of
them because of poor welds. They were returned, fixed, and we
received them back on 19 Jun 97. Once time permits the installation
crew to set the anchors, we will install the stands and the mirror
cans. We received the Brewster window pickoff (one plus a spare)
and are double-checking a few of the critical dimensions before
having them cleaned. Optical cavity components still await final
cleaning to reduce particulates. We are delayed by staff unavailability
due to work on the main accelerator. We determined that a 50/50%
beamsplitter pellicle can be imaged (if sufficiently attenuated)
by our vidicon cameras. We needed to make this determination
before ordering the pellicles that are used for optical cavity
alignment. We measured the sensitivity of an Electrophysics infrared
camera at 633 nm (a possible diagnostic for 3rd harmonic detection.)
We moved the Mach-Zehnder interferometer to the FEL Facility
for testing. It is being used on one optical table until the
other table is aligned. The indium braze used to heat sink a
calcium fluoride window (simulating one of our cavity mirrors)
to a copper holder was cycled to 50 C above ambient nine times
without exhibiting signs of separation. We will now mount a calcium
fluoride mirror (coated for 633 nm) in one of the holders for
mirror heating tests.
We assimilated the available data from the experiments with the
photocathode gun in the ITS, which led to the plans discussed
above in the "Injector Test Stand" section.
There was considerable interplay during the week between our lattice
designer and magnet engineers to assess the relative importance
of various manufacturing complications involving the mounting
of mu-metal sheets in the reverse-bend dipole magnets and the
nature of the corrector magnets in general. The status of these
magnets is summarized above in the "Accelerator Systems"
Investigation of the observed beam scraping at the aperture in
the light box of the photocathode gun with 135 pC bunch charges
gave way to some puzzling results. The simulation (PARMELA) says
that, with a 4 mm laser spot at the cathode, the 4-sigma size
of the beam at that aperture is 1.8 cm, which is a factor of 4/1.8
= 2.22 times smaller than the aperture size there. Furthermore,
with a 6 mm laser spot at the cathode, the 4-sigma size of the
beam at that aperture is 1.5 cm, which is a factor of 4/1.5 =
2.67 times smaller than the aperture size there. So, the code
has thus far offered no explanation about why beam was scraped
at that point. One possibility is that the electrons scraped
away were not from the usable beam itself, but were, for example,
from ghost pulses. We are in process of investigating whether
this was the case.
Finish work continued on the Facility. Outside grading is just
underway to get the grounds restored and the parking areas installed.
The vinyl tile was completed in the break room and upstairs hall.
Ceramic tile work was completed in the bathrooms. The ceramic
fixtures are being installed. The cab and doors are being constructed
on the elevator. The HVAC turns on today (20 Jun 97. On the
Jefferson Lab side we brought 18 workbenches into the optics labs
to support buildup. The first 50 kW RF unit including power supply
and klystron was craned into the Facility and placed. A number
of power supplies and support electronics were also brought in.
Electronic hookup continued in the instrumentation and control
racks. The clean room safety system was completed and is operational.
The cryogenic controls were hooked up to the transfer lines and
the control computer brought on line and connected to the net.
The crane rail for moving the cryounit was qualified early in
the week. The cryounit was extracted from the ITS and installed
in the injector pit. Initial survey and positioning was performed.
Survey and alignment was done on the optical cavity tables.
They are close to final position but will have to await replacement
of a broken positioning bolt on Monday before they are locked
down. A laser interferometer was set on the optical table and
measured maximum vibrational excursions in the tens-of-nanometer
range at 15 to 25 Hz despite lots of mechanical pumps, forklifts,
etc., going in the area. A longer-term drift will be quantified
The clean room is substantially complete and ready for checkout
by Clean Air Technology. The Laser Safety system is also complete
and an electronics rack for the laser electronics has been moved
to the clean room. Installation of the optical bench is awaiting
technician time. The quarter cryounit is installed. Preparations
continue for a cooldown during the July shutdown. Mirror cans
for the drive laser transport are in check. The elbow for the
high voltage power supply was delayed and may not be available
until Aug. 8. Work on installing the power supply has been delayed
due to other conflicting work in the injector area.